OpenStudy (anonymous):

Assuming x > 0, what is the simplified fraction form for

OpenStudy (anonymous):

\[\frac{ 15x \sqrt{3} }{ \sqrt{75x^2} }\]

OpenStudy (raden):

hint : sqrt(75x^2) = sqrt(25 * 3 * x^2) = sqrt(25) * sqrt(3) * sqrt(x^2)